Sketch the orbit described by r = kθ, k is a constant and find the force law for a
central force field that govern this motion.
1
Expert's answer
2019-10-28T11:37:50-0400

We can use the formula for trajectory in the central fields


"\varphi + {\varphi _0} = \int {\frac{{\frac{L}{{{r^2}}}dr}}{{\sqrt {2m(E - U) - \frac{{{L^2}}}{{{r^2}}}} }}}"

Here "E" - the full energy, "L" - angular momentum, "U(r)" - potential energy. The derivation of this equations you may find, for example, in the book "Mechanics: Volume 1 (Course of Theoretical Physics), Landau L.D., Lifshitz E.M.", in paraghaphs 13-14-15.

Now using "r = k\varphi" and letting "{\varphi _0} = 0" we get


"\frac{{\frac{L}{{{r^2}}}dr}}{{\sqrt {2m(E - U) - \frac{{{L^2}}}{{{r^2}}}} }} = {k^{ - 1}}"

Thus


"{k^2}\frac{{{L^2}}}{{{r^4}}} = 2m(E - U) - \frac{{{L^2}}}{{{r^2}}}"

And

"U(r) = E - \frac{1}{{2m}}({k^2}\frac{{{L^2}}}{{{r^4}}} + \frac{{{L^2}}}{{{r^2}}})"

Then the force can be found as "\vec F(r) = - \nabla U" and


"\vec F(r) = - (\frac{{{L^2}}}{m}{r^{ - 3}} + \frac{{2{k^2}{L^2}}}{m}{r^{ - 5}}){{\vec e}_r}"

The trajectory "r = k\varphi" is called Archimedean spiral (at the pircture below we use "k=1")